OpenStudy (anonymous):

find the measure of the numbered angle. a. 230 b. 115 c. 130 d. 125

OpenStudy (anonymous):



OpenStudy (anonymous):

\[ = \angle1 = 1/2(100+130)\] \[=> \angle1 = 230/2 = 115 degrees\]